differential equations

Use the Laplace transform to solve the given initial-value problem. y - sy + 16y = t, Y(0) = 0, y(0) = 1 y(t) =
Answer #1

Given y_8y + lby=t . ylo)=0 y (0)=1. it L {yit) }= y(s) thens L{ y (t)} s. Y(s) - y Co) L{y (+)} = 5+ y(s) – s Y (0) – yIts Y(s)- s(S.35+16) - y(s). its s*(-45-45+16) y (37= tis s($(8-4) - (5-)) y(s) Its Ś (5-4) (5-4) its Ast B = + cstD (5-4) scompare coefficients of son both sides. (A +C) = 0 > -8A + B + D=1 16 A-8B=0 ; .16B = 163 = 1 B=1 16 B = 1 16 16A - 8B= 16A-5y(s) - s +) 5419 () 에 쁩 1% s2 x رای *S 16. visis is (h) s2 32 (-4) Apply inverse laplse Trunstomy. Y () - t 16 eut + 19 teht